OpenStudy (anonymous):

4(2-5x)=6-3(1-3x)? pick a solution: A 29x=17 B 8x=17 C 29x=5 D 8x=5

OpenStudy (anonymous):

8 - 20x = 6 - 3 + 9x 8 - 20x = 3 + 9x 5 = 29x

OpenStudy (anonymous):

OMG! THANX!